# Dynamic-Aware Spatio-temporal Representation Learning for Dynamic MRI Reconstruction

### v1

@@ Abstract
Dynamic MRI reconstruction, one of inverse problems, has seen a surge by the use of deep learning techniques. Especially, the practical difficulty of obtaining ground truth data has led to the emergence of unsupervised learning approaches. A recent promising method among them is implicit neural representation (INR), which defines the data as a continuous function that maps coordinate values to the corresponding signal values. This allows for filling in missing information only with incomplete measurements and solving the inverse problem effectively. Nevertheless, previous works incorporating this method have faced drawbacks such as long optimization time and the need for extensive hyperparameter tuning. To address these issues, we propose Dynamic-Aware INR (DA-INR), an INR-based model for dynamic MRI reconstruction that captures the spatial and temporal continuity of dynamic MRI data in the image domain and explicitly incorporates the temporal redundancy of the data into the model structure. As a result, DA-INR outperforms other models in reconstruction quality even at extreme undersampling ratios while significantly reducing optimization time and requiring minimal hyperparameter tuning.
@@ I INTRODUCTION
  • We show that DA-INR achieves state-of-the-art performance on cardiac cine and DCE liver datasets under various conditions on dynamic MRI reconstruction. It accelerates optimization by 4.63×9.78×4.63$\times$-9.78$\times$4.63 × - 9.78 × and reduces GPU memory usage by 2.15×2.89×2.15$\times$-2.89$\times$2.15 × - 2.89 × compared to the existing methods.
@@ II-A Dynamic MRI reconstruction
Here, $S_{c}\in\mathbb{C}^{(N\times N)\times(N\times N)}$ is a diagonal matrix representing $c^{th}$ coil sensitivity map $(1\leq c\leq C)$, $m_{c}\in\mathbb{C}^{N\times M\times\tau}$ is a $c^{th}$ coil undersampled $(k,t)$ space data, $d\in\mathbb{C}^{N\times N\times\tau}$ is a discrete reconstructed image sequence matrix, $\tau$ is temporal length of image sequence $d$, $N$ is an image size, $M$ is the number of undersampled lines per frame $(N>M)$, and $C$ is the total number of coil channels. $F_{u}\in\mathbb{C}^{(N\times M)\times(N\times N)}$ denotes an undersampled Fourier operator that simulates undersampled data acquisition. If the k-space data is sampled in golden angle radial-way, $F_{u}$ denotes Non-uniform Fast Fourier Transform (NuFFT) operator.
@@ III-A4 Canonical network
$\begin{split}d_{\theta}&=\begin{bmatrix}f_{\theta}(1,1,t)&\cdots&f_{\theta}(1,rW,t)\\ \vdots&\cdots&\vdots\\ f_{\theta}(rH,1,t)&\cdots&f_{\theta}(rH,rW,t)\\ \end{bmatrix},\end{split}$ (7)
@@ III-A6 Optimization
This equation is based on Eq. 4. $\mathbf{R}(d_{\theta})$ is composed of temporal TV and low rank regularization terms. $TV_{t}(\cdot)$ denotes a temporal TV operator and $||\cdot||_{*}$ denotes a sequent operator of singular value decomposition and sum of singular values. $\lambda_{S}$ and $\lambda_{L}$ are the weights for the temporal TV and the low rank regularization, respectively. In contrast, DA-INR fits all the frames to the shared canonical space, constraining the solution space through the canonical network $\Psi_{x}$. The fitting process in DA-INR can be interpreted as mapping $(x,y)$ in the canonical space to the corresponding intensity values, while other frames retrieve these signal values at their cell points based on the deformation field $[\Delta x_{k}]_{k=1}^{N}$, which accounts for the temporal difference (Fig.2 (b)). Since the canonical space is shared across all frames, this design induces interconnection within the image sequence, allowing the model to focus on signals in the canonical space and enabling faster compensation for artifacts caused by undersampling in the other frames.
@@ IV Experiments
Figure 3: Sensitivity analysis of Feng et al. [12] on cardiac cine data. “R2” denotes the use of relative L2 loss [18] for data-consistency optimization, while “L1” refers to the use of L1 loss. $\lambda_{L}$ and $\lambda_{S}$ are weights for low rank and temporal TV regularization, respectively. Lastly, “$step>20$” denotes that temporal TV regularization is turned on after training step $20$.
@@ IV-B1 Retrospective cardiac cine data
Cardiac cine data was obtained using a 3T whole-body MRI scanner (Siemens Tim Trio) equipped with a 32-element cardiac coil array. The acquisition utilized a bSSFP sequence with prospective cardiac gating. The imaging parameters were as follows: FOV $=(300\times 300)$ mm2, acquisition matrix $=(128\times 128)$, TE/TR=$1.37/2.7$ ms, receiver bandwidth = 1184 Hz/pixel, and flip angle $=40^{\circ}$. The number of frames was 23 and the temporal resolution was 43.2 ms. The full-sampled k-space data is used as ground truth (GT). To simulate a retrospective undersampling pattern, we adopt a 2D golden-angle radial acquisition scheme, where the spokes repeatedly traverse the center of k-space, rotating with a step of $111.25^{\circ}$. It is applied to ground truth with multi-coil NuFFT to obtain the undersampled radial trajectories of Fibonacci numbers [19].
@@ IV-B2 Dynamic Contrast-Enhanced (DCE) liver data
The DCE liver MRI scan was conducted on a healthy volunteer using axial orientation and breath-holding techniques by a whole-body 3 Tesla MRI system (MAGNETOM Verio/Avanto, Siemens AG, Erlangen, Germany), employing a combination of body-matrix and 12-element spine coil array. For data acquisition, a radial stack-of-stars 3D Fast Low Angle Shot (FLASH) pulse sequence with golden-angle ordering was utilized. The imaging parameters were as follows: FOV = $370\times 370$ mm2, TR/TE=3.83 ms/1.71 ms, acquisition matrix = $384\times 384$, total spoke for one slice $=600$. In our experiments, 34 consecutive spokes are used for each frame, which provides 17 temporal frames. The reconstructed image matrix for each frame is 384$\times$384. The acceleration factor $(AF)$ is $384/34=11.3$.
@@ IV-D Implementation details
[…]
Figure 4: Visual comparisons between results of $AF=9.8$ in cardiac cine data reconstruction at diastole and systole. The upper row is the reconstruction output in the $(y-x)$ domain for each method and the below row is the absolute error map between ground truth and the reconstructed output of each method. PSNR values are specific to each frame.
Figure 5: Visual comparisons between results of $AF=25.6$ in cardiac cine data reconstruction at diastole and systole. The upper row is the reconstruction output in the $(y-x)$ domain for each method and the below row is the absolute error map between ground truth and the reconstructed output of each method. PSNR values are specific to each frame.
@@ V-A Retrospective cardiac cine data
Fig. 4 and Fig. 5 present the visual comparisons between our method and the existing methods for cardiac cine data reconstruction at $AF=9.8$ and $AF=25.6$, respectively. We evaluate the reconstructed frames for each method during the diastolic and systolic phases. When undersampling is performed with only five spokes per frame ($AF=25.6$), NuFFT struggles to accurately capture the cardiac structure. For HashINR, reconstruction quality improves as the number of sampled spokes increases; however, residual noise persists even at $AF=9.8$. While GRASP sometimes achieves high PSNR, it struggles to reconstruct fine structural details, such as the shape of the papillary muscle, under both undersampling ratios. TD-DIP shows limitations in accurately capturing the precise contraction of the papillary muscle during the systolic phase at $AF=25.6$. In contrast, our method closely approximates the ground truth, achieving high fidelity in both the diastolic and systolic phases. The quantitative error—computed as the sum of the squared differences between the reconstructed and ground truth images—is the smallest among all methods. Fig. 6 presents the qualitative reconstruction results for cardiac cine data in the $(x-t)$ domain. While TD-DIP and GRASP produce satisfactory results in the $(y-x)$ domain, their performance deteriorates in the $(x-t)$ domain. HashINR achieves reconstructions in the $(x-t)$ domain comparable to the ground truth at $AF=9.8$, though with some noise, and suffers from pronounced blurring along the time dimension at $AF=25.6$. In contrast, our method demonstrates superior reconstruction quality, with reduced noise and minimal blurring.
Figure 6: Qualitative results of cardiac cine data reconstruction in the $(x-t)$ domain. A white line at the heart region indicates the cross section that is visualized.
@@ V-B Dynamic Contrast-Enhanced (DCE) liver data
Fig. 7 presents the qualitative reconstruction results and the corresponding signal intensity flowmap for DCE liver data reconstruction performed with an undersampling ratio of 34 spokes per frame $(AF=11.3)$. GRASP reconstructs images in the $(y-x)$ domain with minimal noise, as shown (Fig. 7 (a)). However, it suffers from low temporal fidelity, as evident in the signal intensity flowmap (Fig. 7 (b)). On the other hand, HashINR exhibits good temporal fidelity in the flowmap, but its outputs display noticeable noise in the $(y-x)$ domain, overfitting to the undersampled frames. TD-DIP produces the reconstructions characterized by overly smooth appearances. This smoothness results in a failure to accurately delineate fine structural and contrast changes (Fig. 7 (a)). Consequently, it achieves the lowest temporal fidelity among the compared approaches, as reflected in the signal intensity flowmap in Fig. 7 (b). In contrast, our proposed method achieves a prominent performance, delivering both high temporal fidelity and phase-specific contrast changes. The reconstructions in Fig. 7 (a) are well-defined, showing clear structural details and accurate phase enhancements. Furthermore, our signal intensity flowmap in Fig. 7 (b) shows a strong capacity of our model to preserve temporal dynamics, capturing the changes in signal intensity over time with high accuracy.
@@ V-C Time consumption and GPU memory usage
Figure 8: Qualitative results in spatial interpolation (×1.2,×1.5,×2)($\times$ 1.2$\times$ 1.5$\times$ 2)( × 1.2 , × 1.5 , × 2 ) task at $AF=9.8$. PSNR values are specific to each frame.
Figure 9: Comprehensive results in temporal interpolation (×2,×3)($\times$ 2$\times$ 3)( × 2 , × 3 ) task. (a), (b) Temporal interpolation results at $AF=6.1$ for (×2)($\times$ 2)( × 2 ) and (×3)($\times$ 3)( × 3 ), respectively. PSNR and SSIM values are specific to each frame. (c) The visual comparison of each method in the $(x-t)$ domain. A white line at the heart region of GT $(y-x)$ image indicates the cross section that is visualized.
@@ V-D Results of spatial and temporal interpolation
We conduct experiments of spatial interpolation (×1.2,×1.5,×2)($\times$ 1.2$\times$ 1.5$\times$ 2)( × 1.2 , × 1.5 , × 2 ) and temporal interpolation (×2,×3)($\times$ 2$\times$ 3)( × 2 , × 3 ) to validate the internal continuity of our model. For spatial interpolation, we employ bilinear interpolation, HashINR, and our proposed method. For temporal interpolation, we utilize HashINR, TD-DIP, and our method.
@@
For temporal interpolation, we optimize the models by taking as inputs spatio-temporal coordinates corresponding to evenly subsampled frames according to the sampling ratio (×2)($\times$ 2)( × 2 ) or (×3)($\times$ 3)( × 3 ). During inference, the coordinates from all frames are fed into the models to generate whole image sequence. We extract the image features of two neighboring frames of time $t$ because we can not exploit the image features corresponding to the unseen time coordinates during optimization. Fig. 9 (a) and (b) show the overall results of reconstruction with an undersampling ratio of 21 spokes sampled per frame $(AF=6.1)$ and Fig. 9 (c) visualizes output of each method in the $(x-t)$ dimension. As frame number goes from $7$ to $9$, the papillary muscle detaches from the wall in ground truth. The output of HashINR at time $9$ shows a clear detachment of the muscle from the wall, but the interpolated output at time $8$ completely loses structural integrity and its visual quality is extremely poor. As shown in Fig. 9 (c), there are temporal vacancy artifacts in the outputs of HashINR in the $(x-t)$ domain. TD-DIP can not reconstruct the definite anatomy of the muscle and its contraction through the time. As a result, the outputs of TD-DIP are blurry and indistinct in the $(x-t)$ domain, as shown in Fig. 9 (c). However, our results show a distinct detachment of the muscle from the wall with the best quantitative and qualitative results (Fig. 9 (a) and (b)). In addition, ours show the highest fidelity in the $(x-t)$ domain with less noise or artifacts with optimistic interpolation capacity (Fig. 9 (c)).

### v2

@@ Abstract
Dynamic MRI reconstruction, one of inverse problems, has seen a surge by the use of deep learning techniques. Especially, the practical difficulty of obtaining ground truth data has led to the emergence of unsupervised learning approaches. A recent promising method among them is implicit neural representation (INR), which defines the data as a continuous function that maps coordinate values to the corresponding signal values. This allows for filling in missing information only with incomplete measurements and solving the inverse problem effectively. Nevertheless, previous works incorporating this method have faced drawbacks such as long optimization time and the need for extensive hyperparameter tuning. To address these issues, we propose Dynamic-Aware INR (DA-INR), an INR-based model for dynamic MRI reconstruction that captures the spatial and temporal continuity of dynamic MRI data in the image domain and explicitly incorporates the temporal redundancy of the data into the model structure. As a result, DA-INR outperforms other models in reconstruction quality even at extreme undersampling ratios while significantly reducing optimization time and requiring minimal hyperparameter tuning. Our code is available at here.
@@ I INTRODUCTION
  • We show that DA-INR achieves state-of-the-art performance on cardiac cine and DCE liver datasets under various conditions on dynamic MRI reconstruction. It accelerates optimization by $4.63\times-9.78\times$ and reduces GPU memory usage by $2.15\times-2.89\times$ compared to the existing methods.
@@ II-A Dynamic MRI reconstruction
Here, $S_{c}\in\mathbb{C}^{(N\times N)\times(N\times N)}$ is a diagonal matrix representing $c^{th}$ coil sensitivity map $(1\leq c\leq C)$, $m_{c}\in\mathbb{C}^{N\times M\times\tau}$ is a $c^{th}$ coil undersampled $(k,t)$ space data, $d\in\mathbb{C}^{N\times N\times\tau}$ is a discrete reconstructed image sequence matrix, $\tau$ is temporal length of image sequence $d$, NN is an image size, $M$ is the number of undersampled lines per frame $(N>M)$, and $C$ is the total number of coil channels. $F_{u}\in\mathbb{C}^{(N\times M)\times(N\times N)}$ denotes an undersampled Fourier operator that simulates undersampled data acquisition. If the k-space data is sampled in golden angle radial-way, $F_{u}$ denotes Non-uniform Fast Fourier Transform (NuFFT) operator.
@@ III-A4 Canonical network
where $Re$ and $Im$ denote real and imaginary components of the cell position $\mathbf{x}^{\prime}$ in the complex-valued image, respectively, and $\langle\cdot,\cdot\rangle$ means channel-wise concatenation. The final output image $d_{\theta}\in\mathbb{C}^{rH\times rW}$ at time $t$ is defined as:
@@ III-A6 Optimization
This equation is based on Eq. 4. $\mathbf{R}(d_{\theta})$ is composed of temporal TV and low rank regularization terms. $TV_{t}(\cdot)$ denotes a temporal TV operator and $||\cdot||_{*}$ denotes a sequent operator of singular value decomposition and sum of singular values. $\lambda_{S}$ and $\lambda_{L}$ are the weights for the temporal TV and the low rank regularization, respectively. In contrast, DA-INR fits all the frames to the shared canonical space, constraining the solution space through the canonical network $\Psi_{x}$. The fitting process in DA-INR can be interpreted as mapping $(x,y)$ in the canonical space to the corresponding intensity values, while other frames retrieve these signal values at their cell points based on the deformation field $\Delta x_{k}]_{k=1}^{N}$, which accounts for the temporal difference (Fig.2 (b)). Since the canonical space is shared across all frames, this design induces interconnection within the image sequence, allowing the model to focus on signals in the canonical space and enabling faster compensation for artifacts caused by undersampling in the other frames.
@@ IV Experiments
Figure 3: Sensitivity analysis of Feng et al. [12] on cardiac cine data. “R2” denotes the use of relative L2 loss [18] for data-consistency optimization, while “L1” refers to the use of L1 loss. $\lambda_{L}$ and $\lambda_{S}$ are weights for low rank and temporal TV regularization, respectively. Lastly, “step>20step>20” denotes that temporal TV regularization is turned on after training step 2020.
@@ IV-B1 Retrospective cardiac cine data
Cardiac cine data was obtained using a 3T whole-body MRI scanner (Siemens Tim Trio) equipped with a 32-element cardiac coil array. The acquisition utilized a bSSFP sequence with prospective cardiac gating. The imaging parameters were as follows: FOV $(300\times 300)$ mm2, acquisition matrix $(128\times 128)$, TE/TR=$1.37/2.7$ ms, receiver bandwidth = 1184 Hz/pixel, and flip angle $40^{\circ}$. The number of frames was 23 and the temporal resolution was 43.2 ms. The full-sampled k-space data is used as ground truth (GT). To simulate a retrospective undersampling pattern, we adopt a 2D golden-angle radial acquisition scheme, where the spokes repeatedly traverse the center of k-space, rotating with a step of $111.25^{\circ}$. It is applied to ground truth with multi-coil NuFFT to obtain the undersampled radial trajectories of Fibonacci numbers [19].
@@ IV-B2 Dynamic Contrast-Enhanced (DCE) liver data
The DCE liver MRI scan was conducted on a healthy volunteer using axial orientation and breath-holding techniques by a whole-body 3 Tesla MRI system (MAGNETOM Verio/Avanto, Siemens AG, Erlangen, Germany), employing a combination of body-matrix and 12-element spine coil array. For data acquisition, a radial stack-of-stars 3D Fast Low Angle Shot (FLASH) pulse sequence with golden-angle ordering was utilized. The imaging parameters were as follows: FOV = $370\times 370$ mm2, TR/TE=3.83 ms/1.71 ms, acquisition matrix = $384\times 384$, total spoke for one slice =600=600. In our experiments, 34 consecutive spokes are used for each frame, which provides 17 temporal frames. The reconstructed image matrix for each frame is $\times384$. The acceleration factor $(AF)$ is $384/34=11.3$.
@@ IV-D Implementation details
For the deformation network and the canonical network, we use a small MLP with five hidden layers, each comprising 64 neurons, followed by a ReLU activation function. The MLP of the canonical network predicts two channels representing the real and imaginary components of the complex-valued MRI images, with no activation function applied to the final layer. During the optimization process, all spatio-temporal coordinates are assembled in a single batch and normalized to the range [1,1][-1,1]. The batch size is set to 1. We use the AdamW [20] optimizer with a learning rate of 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$ during optimization. The non-Cartesian Fourier undersampling operation is executed using the NuFFT package1, facilitating rapid computation and gradient backpropagation on a GPU. Feng et al. [12] tunes the hyperparameters $L,T,F,N_{min}$, and $b$ of its hash encoder according to each data type and AF, while we use consistent values across all data types and AFs as $L=16,T=2^{19},F=2,N_{min}=16$, and $b=2$.
Figure 4: Visual comparisons between results of $AF=9.8$ in cardiac cine data reconstruction at diastole and systole. The upper row is the reconstruction output in the (yx)(y-x) domain for each method and the below row is the absolute error map between ground truth and the reconstructed output of each method. PSNR values are specific to each frame.
Figure 5: Visual comparisons between results of $AF=25.6$ in cardiac cine data reconstruction at diastole and systole. The upper row is the reconstruction output in the (yx)(y-x) domain for each method and the below row is the absolute error map between ground truth and the reconstructed output of each method. PSNR values are specific to each frame.
@@ V-A Retrospective cardiac cine data
Fig. 4 and Fig. 5 present the visual comparisons between our method and the existing methods for cardiac cine data reconstruction at $AF=9.8$ and $AF=25.6$, respectively. We evaluate the reconstructed frames for each method during the diastolic and systolic phases. When undersampling is performed with only five spokes per frame ($AF=25.6$), NuFFT struggles to accurately capture the cardiac structure. For HashINR, reconstruction quality improves as the number of sampled spokes increases; however, residual noise persists even at $AF=9.8$. While GRASP sometimes achieves high PSNR, it struggles to reconstruct fine structural details, such as the shape of the papillary muscle, under both undersampling ratios. TD-DIP shows limitations in accurately capturing the precise contraction of the papillary muscle during the systolic phase at $AF=25.6$. In contrast, our method closely approximates the ground truth, achieving high fidelity in both the diastolic and systolic phases. The quantitative error—computed as the sum of the squared differences between the reconstructed and ground truth images—is the smallest among all methods. Fig. 6 presents the qualitative reconstruction results for cardiac cine data in the (xt)(x-t) domain. While TD-DIP and GRASP produce satisfactory results in the (yx)(y-x) domain, their performance deteriorates in the (xt)(x-t) domain. HashINR achieves reconstructions in the (xt)(x-t) domain comparable to the ground truth at $AF=9.8$, though with some noise, and suffers from pronounced blurring along the time dimension at $AF=25.6$. In contrast, our method demonstrates superior reconstruction quality, with reduced noise and minimal blurring.
Figure 6: Qualitative results of cardiac cine data reconstruction in the (xt)(x-t) domain. A white line at the heart region indicates the cross section that is visualized.
@@ V-B Dynamic Contrast-Enhanced (DCE) liver data
Fig. 7 presents the qualitative reconstruction results and the corresponding signal intensity flowmap for DCE liver data reconstruction performed with an undersampling ratio of 34 spokes per frame $(AF=11.3)$. GRASP reconstructs images in the (yx)(y-x) domain with minimal noise, as shown (Fig. 7 (a)). However, it suffers from low temporal fidelity, as evident in the signal intensity flowmap (Fig. 7 (b)). On the other hand, HashINR exhibits good temporal fidelity in the flowmap, but its outputs display noticeable noise in the (yx)(y-x) domain, overfitting to the undersampled frames. TD-DIP produces the reconstructions characterized by overly smooth appearances. This smoothness results in a failure to accurately delineate fine structural and contrast changes (Fig. 7 (a)). Consequently, it achieves the lowest temporal fidelity among the compared approaches, as reflected in the signal intensity flowmap in Fig. 7 (b). In contrast, our proposed method achieves a prominent performance, delivering both high temporal fidelity and phase-specific contrast changes. The reconstructions in Fig. 7 (a) are well-defined, showing clear structural details and accurate phase enhancements. Furthermore, our signal intensity flowmap in Fig. 7 (b) shows a strong capacity of our model to preserve temporal dynamics, capturing the changes in signal intensity over time with high accuracy.
@@ V-C Time consumption and GPU memory usage
Figure 8: Qualitative results in spatial interpolation $(\times 1.2,\times 1.5,\times 2)$ task at $AF=9.8$. PSNR values are specific to each frame.
Figure 9: Comprehensive results in temporal interpolation $(\times 2,\times 3)$ task. (a), (b) Temporal interpolation results at $AF=6.1$ for $(\times 2)$ and $(\times 3)$, respectively. PSNR and SSIM values are specific to each frame. (c) The visual comparison of each method in the (xt)(x-t) domain. A white line at the heart region of GT (yx)(y-x) image indicates the cross section that is visualized.
@@ V-D Results of spatial and temporal interpolation
We conduct experiments of spatial interpolation $(\times 1.2,\times 1.5,\times 2)$ and temporal interpolation $(\times 2,\times 3)$ to validate the internal continuity of our model. For spatial interpolation, we employ bilinear interpolation, HashINR, and our proposed method. For temporal interpolation, we utilize HashINR, TD-DIP, and our method.
@@
For temporal interpolation, we optimize the models by taking as inputs spatio-temporal coordinates corresponding to evenly subsampled frames according to the sampling ratio $(\times 2)$ or $(\times 3)$. During inference, the coordinates from all frames are fed into the models to generate whole image sequence. We extract the image features of two neighboring frames of time $t$ because we can not exploit the image features corresponding to the unseen time coordinates during optimization. Fig. 9 (a) and (b) show the overall results of reconstruction with an undersampling ratio of 21 spokes sampled per frame $(AF=6.1)$ and Fig. 9 (c) visualizes output of each method in the (xt)(x-t) dimension. As frame number goes from $7$ to $9$, the papillary muscle detaches from the wall in ground truth. The output of HashINR at time $9$ shows a clear detachment of the muscle from the wall, but the interpolated output at time $8$ completely loses structural integrity and its visual quality is extremely poor. As shown in Fig. 9 (c), there are temporal vacancy artifacts in the outputs of HashINR in the (xt)(x-t) domain. TD-DIP can not reconstruct the definite anatomy of the muscle and its contraction through the time. As a result, the outputs of TD-DIP are blurry and indistinct in the (xt)(x-t) domain, as shown in Fig. 9 (c). However, our results show a distinct detachment of the muscle from the wall with the best quantitative and qualitative results (Fig. 9 (a) and (b)). In addition, ours show the highest fidelity in the (xt)(x-t) domain with less noise or artifacts with optimistic interpolation capacity (Fig. 9 (c)).